AT 9:00 A.M. A TRUCK LEAVES THE TRUCK YARD AND
TRAVELS WEST AT A RATE OF 35 MI/HR.
TWO HOURS LATER, A SECOND TRUCK LEAVES ALONG THE SAME ROUTE
TRAVELING AT 70 MI/HR.
WHEN WILL THE SECOND TRUCK CATCH UP TO THE FIRST TRUCK?
THANKS
JMS

Answer by checkley71(8403)

DISTANCE=RATE*TIME
SEEING AS THE DISTANCES ARE THE SAME WE HAVE:
35T=70(T-2)
35T=70T-140
35T-70T=-140
-35T=-140
T=-140/-35
T=4 HOURS THEY WILL MEET.
PROOF
35*4=70(4-2)
140=70*2
140=140